out of 41 a al l Δ Download Answer ALL questions 4. This is a simplified bank balance sheet. Suppose the bank is required to maintain a cash reserve at 10% of total liabilities. Now a large depositor withdraws £40M. i. Does the banks balance sheet meet the cash reserve requirement after the ii. How much new deposits are needed at the minimum in order to restore the withdrawal? cash reserve to the required level? Liabilities & Capital (EMillions) Assets (EMillions) 500 Cash reserves75 Marketable 100 securities Loans Capital 100 425 10 Marks

question 4 please
Answer #1

4 Cash Reserves = 10% of Liabilities
Total Liabilities = 600 million
Withdrawal will reduce deposits in liabilities  side and cash in assets side.

i) So Total liabilities = 600-40 = 560
Total Cash = 75-40 = 35
Reserve Requirement = 10%
So 10% of 560 = 560*10% = 56
Hence the cash reserve requirement is not met.

ii) Let New deposits required be x
So cash reserve requirement with new deposit = (x+35)/(x+560) = 10%
x + 35 = 0.1 x + 56
0.9x = 21
x = 21/0.9 = 23.33 million

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well